OpenStudy (anonymous):

I need help on how to solve these. Jennifer swims 4 complete lengths of a 50m pool. a. What distance did she swim? b. What is her displacement?

OpenStudy (astrophysics):

Note that distance is a scalar quantity and displacement is a vector quantity, that means distance is just a magnitude you would have to add all the times Jennifer swam to get the total distance, but the displacement is a bit more tricky. It depends on the direction, so say we have a square  the distance here would be (20+20+20+20)m = 80m, but the displacement, if we say you go all around, then you would do up +down, and left +right, denoting up and right as positive we would have 20m+(-20m) = 0 and left + right would be -20m+20m = 0 so then our displacement would be 0. Hope that helps!